OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). a = 3b - 4 a + b =16

OpenStudy (anonymous):

this equation can be solved because from the first equation you have a=3b-4 so if we substitute that in the second equation for a, we can then solve for b. doing so gives us this: 3b-4+b=16 4b-4=16 4b=20 b=5 now that we know b, we can plug that back into either original equation and solve for a: a=3(5)-4 a=15-4 a=11 so your ordered pair will look like this: (5,11)